Required information Exercise 4-16 (Algo) Working with a Segmented Income Statement; Break-Even Analysis [LO4-4, LO4-5] [The following information applies to the questions displayed below.] Raner, Harris & Chan is a consulting firm that specializes in information systems for medical and dental clinics. The firm has two offices—one in Chicago and one in Minneapolis. The firm classifies the direct costs of consulting jobs as variable costs. A contribution format segmented income statement for the company’s most recent year is given: Office Total Company Chicago Minneapolis Sales $ 1,012,500 100.0 % $ 202,500 100 % $ 810,000 100 % Variable expenses 546,750 54.0 % 60,750 30 % 486,000 60 % Contribution margin 465,750 46.0 % 141,750 70 % 324,000 40 % Traceable fixed expenses 226,800 22.4 % 105,300 52 % 121,500 15 % Office segment margin 238,950 23.6 % $ 36,450 18 % $ 202,500 25 % Common fixed expenses not traceable to offices 162,000 16.0 % Net operating

income $ 76,950 7.6 % Exercise 4-16 (Algo) Part 1 Required: 1-a. Compute the companywide break-even point in dollar sales. 1-b. Compute the break-even point for the Chicago office and for the Minneapolis office. 1-c. Is the companywide break-even point greater than, less than, or equal to the sum of the Chicago and Minneapolis break-even points?

2. By how much would the company’s net operating income increase if Minneapolis increased its sales by $101,250 per year? Assume no change in cost behavior patterns.

3. Assume that sales in Chicago increase by $67,500 next year and that sales in Minneapolis remain unchanged. Assume no change in fixed costs.

a. Prepare a new segmented income statement for the company. (Round your percentage answers to 1 decimal place (i.e. 0.1234 should be entered as 12.3).)
Answer #1
Break-even point = Fixed expenses/Contribution margin ratio
1a
Break-even point in dollar sales 845217 =(226800+162000)/46%
1b
Break even point
Chicago office 150429 =105300/70%
Minneapolis office 303750 =121500/40%
1c
Greater than
2
Increase in sales 101250
X CM ratio 40%
Net operating income increase 40500
3
Total company Chicago Minneapolis
Amount % Amount % Amount %
Sales 1080000 100.0% 270000 100.0% 810000 100.0%
Variable expenses 567000 52.5% 81000 30.0% 486000 60.0%
Contribution margin 513000 47.5% 189000 70.0% 324000 40.0%
Traceable fixed expenses 226800 21.0% 105300 39.0% 121500 15.0%
Office segment margin 286200 26.5% 83700 31.0% 202500 25.0%
Common fixed expenses not traceable 162000 15.0%
Net operating income 124200 11.5%